The students of same age group from two different schools were compared for variability in their mathematical skill. A random sample of 25 pupils from one school had a variance of 16 marks while a random sample of 22 pupils from the other school had a variance of 8 marks. Examine if




the difference in variability is significant IF at (24.21) at 5% level of significance

1
Expert's answer
2021-12-14T14:31:34-0500

F test for variances, using F distribution (dfnum=24,dfdenom=21) (two-tailed) 


"H_0:\sigma_1=\sigma_2" , there is no difference in variability in two groups

"H_a:\sigma_1\neq\sigma_2" , there is difference in variability in two groups


"F=\frac{\sigma_1^2}{\sigma_2^2}=\frac{16}{8}=2"


critical value:

"F_{crit}=2.05"


Since "F<F_{crit}" we accept null hypothesis. There is no difference in variability in two groups.